# X. Wang1,2,+    Siming Fu1,+    Qihan Huang1,2    Wanggui He1    Hao Jiang1,† 1

### v1

@@ Abstract
Recent advancements in text-to-image generation models have dramatically enhanced the generation of photorealistic images from textual prompts, leading to an increased interest in personalized text-to-image applications, particularly in multi-subject scenarios. However, these advances are hindered by two main challenges: firstly, the need to accurately maintain the details of each referenced subject in accordance with the textual descriptions; and secondly, the difficulty in achieving a cohesive representation of multiple subjects in a single image without introducing inconsistencies. To address these concerns, our research introduces the MS-Diffusion framework for layout-guided zero-shot image personalization with multi-subjects. This innovative approach integrates grounding tokens with the feature resampler to maintain detail fidelity among subjects. With the layout guidance, MS-Diffusion further improves the cross-attention to adapt to the multi-subject inputs, ensuring that each subject condition acts on specific areas. The proposed multi-subject cross-attention orchestrates harmonious inter-subject compositions while preserving the control of texts. Comprehensive quantitative and qualitative experiments affirm that this method surpasses existing models in both image and text fidelity, promoting the development of personalized text-to-image generation.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion methodologies [27, 29, 3] have propelled the field to new heights, realizing unprecedented levels of photorealism while demonstrating a refined ability to conform to textual prompts. These achievements have spurred the development of a broad spectrum of applications, most prominently in the domain of personalized T2I (P-T2I) models, which are tasked with the complex undertaking of assimilating and regenerating novel visual concepts or subjects across diverse contexts with a heightened demand for conceptual and compositional fidelity. Despite fine-tuning-based techniques such as DreamBooth [28] and Textual Inversion [9] yield results with considerable accuracy, they necessitate extensive resources for tuning individual instances and for the storage of multiple models, which renders them less feasible for widespread application. To circumvent these resource-intensive requirements, fine-tuning-free alternatives have come to the fore.
Single-subject driven personalization methods, IP-Adapter [38] and ELITE [36] for instance, introduce a specialized cross-attention mechanism that distinctly processes text and image features, thereby affording the possibility to employ reference images directly as visual prompts within the model. Furthermore, recent works have employed multi-subject driven customization methodologies to concatenate visual concepts with textual prompts, offering a glimpse of the potential in techniques like SSR-Encoder [40], $\lambda$-ECLIPSE [20], Emu2 [30], and KOSMOS-G [19]. These models harness identity data and amalgamate it with text via cross-attention, exhibiting proficiency in adjusting textures. Nevertheless, above zero-shot personalization methods encounter limitations, notably in adapting a pressing question pertains to the congruence of granular details between the subject depicted in the synthesized imagery and its corresponding subjects, along with the degree of semblance between the content of the generated image and associated textual descriptions. The challenge is further amplified in scenarios requiring the personalization of multiple subjects. Especially, the challenge of ensuring harmonious representation when multiple subjects are incorporated—specifically, the elucidation of whether the resultant image manifests any discordant elements or deleterious interactions in accordance with textual directives and multi-subject referential controls. As illustrated in Figure 2, multi-subject personalization methods frequently incur notable detail inaccuracies in a fine-tuning-free framework and often lead to subject neglect, subject overcontrol, and subject-subject conflict issue within the generated images.
@@
Textual Inversion [9]
DreamBooth [28]
@@
BLIP-Diffusion [15]
IP-Adapter [38]
@@
Kosmos-G [19]
$\lambda$-ECLIPSE [20]
SSR-Encoder [40]
@@ 2.1 Text-to-image Generation
Text-to-image generative models [27, 29, 3, 2] are capable of producing high-quality images using user-provided text prompts. In recent times, diffusion-based models have shown strong performance in text-to-image tasks. Stable Diffusion [27] proposes conducting the diffusion process in latent space rather than pixel space, which reduces the sampling steps without compromising image quality. Kandinsky [26] takes both text embedding and image embedding as conditions to generate images more controllablely. DALLE-3 [3] recaptions the training data pairs and utilizes T5-XXL [7] as the text encoder to strengthen the prompt-following ability. StableCascade [22] presents a cascaded architecture to leverage outputs of front stages as priors, further reducing the latent space. PixArt-$\alpha$ [5] also employs a large T5 text encoder and replaces the original U-Net backbone with a transformer [21]. These models focus on the basic text-to-image ability and cannot handle the situation when users provide specific subjects.
To finely control text-to-image generation, some diffusion models [41, 17, 35] support users in providing layout guidance. Layout Diffusion [41] and GLIGEN [17] input positions and labels of bounding boxes into the diffusion model and train it to learn the layout information. DenseDiffusion [12] develops a training-free method and modulates the attention maps in the inference phase. Instance Diffusion [34] and MIGC [42] extend the layout-conditioned diffusion to the instance level, enabling the model to generate multiple objects with precise quantities. While layout-guided diffusion models have robust controllability, they cannot reference specific concepts, which is emphasized in personalized text-to-image generation.
@@ 2.2 Text-to-image Personalization
Text-to-image personalization [28, 9, 14, 10, 24] has attracted much attention in the community for its powerful referencing capability of both text and image prompts. Textual inversion [9] and DreamBooth [28] utilize an identifier in the text to bind the visual concept through fine-tuning. IP-Adapter [38] proposes a zero-shot personalized model by projecting the image embedding to the cross-attention layers. InstanceID [33] develops an approach for identity personalization, replacing the image encoder with a face encoder and employing ControlNet [39] to integrate the face landmarks. To narrow the gap between image and text prompts, Kosmos-G [19] and $\lambda$-ECLIPSE [20] conduct multi-modal training to unite the inputs by text-image interleaving. SSR-Encoder [40] design a query network to extract a single subject from images with multiple subjects for personalization.
Though past research in this field has significantly enhanced the ability to reference single subjects, few zero-shot multi-subject personalized models. Moreover, existing related works struggle to address conflicts in personalized generation with multiple subjects and generate bad results, which is precisely the focus of our work, to discuss and resolve these issues.
Figure 3: The overall pipeline of MS-Diffusion. It introduces two pivotal enhancements to the model: the grounding resampler and multi-subject cross-attention mechanisms. Firstly, the grounding resampler adeptly assimilates visual information, correlating it with specific entities and spatial constraints. Subsequently, a targeted cross-attention mechanism facilitates precise interactions between the image condition and the diffusion latent within the multi-subject attention layers. Throughout the training phase, all components of the pre-existing diffusion model remain frozen.
@@ 3.1 Preliminaries
As a widely used diffusion model, Stable Diffusion (SD) [27] conducts the diffusion process in the latent space. Given an image and a text prompt, SD encodes them into latent code $\mathbf{z}$ and condition embedding $\mathbf{c_{t}}$ utilizing VAE [32] and CLIP [25] text encoder, respectively. In zero-shot image personalization architectures like IP-adapter [38], images can also be considered a condition of the diffusion model. Specifically, a subject image is encoded to image embeddings by an image encoder and then projected into the original condition space of the diffusion model denoted as $\mathbf{c}_{i}$. For a timestep $t$ which is uniformly sampled from a fixed range, the model $\theta$ predicts the noise $\boldsymbol{\epsilon}_{\theta}$ and is optimized through the objective:
$\mathcal{L}_{IP}=\mathbb{E}_{\mathbf{z},\mathbf{c},\epsilon,t}\left[\left\|{\epsilon-\boldsymbol{\epsilon}}_{\theta}\left(\mathbf{z}_{t}\mid\mathbf{c_{t}},\mathbf{c_{i}},t\right)\right\|_{2}^{2}\right]$ (1)
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In this work, we employ $\mathcal{L}_{IP}$ with SDXL [23] as the pre-trained model, which contains two CLIP text encoders and additional condition inputs besides $\mathbf{c}$ and $t$, omitted for brevity.
@@
In SD [27], both $\mathbf{c}_{i}$ and $\mathbf{c}_{t}$ are integrated into the U-Net backbone through cross-attention layers:
$\text{Attn}\left(\mathbf{Q},\mathbf{K}_{i},\mathbf{K}_{t},\mathbf{V}_{i},\mathbf{V}_{t}\right)=\gamma\cdot\mathbf{z}_{img}+\mathbf{z}_{txt}=\gamma\cdot\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}_{i}}{\sqrt{d}}\right)\mathbf{V}_{i}+\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}_{t}}{\sqrt{d}}\right)\mathbf{V}_{t}$ (2)
where $\mathbf{Q}=\mathbf{z}W_{q}$, $\mathbf{K_{i}}=\mathbf{c}_{i}W_{k}^{i}$, $\mathbf{K_{t}}=\mathbf{c}_{t}W_{k}^{t}$, $\mathbf{V_{i}}=\mathbf{c}_{i}W_{v}^{i}$, $\mathbf{V_{t}}=\mathbf{c}_{t}W_{v}^{t}$, and $W_{q}$, $W_{k}$, $W_{v}$ are corresponding projection weight matrices. And $d$ represents the dimensionality of the key vectors Note that the key and value matrix of $\mathbf{c}_{i}$ and $\mathbf{c}_{t}$ are independent of each other to decouple conditions of different modalities. Previous studies [11, 31] have found that attention maps $\mathbf{A}=\text{Softmax}\left(\mathbf{Q}\mathbf{K}/\sqrt{d}\right)$ can reflect the attribution relation between generated images and conditions, which means that they determine the effect of condition controls.
@@ 3.2 Data Construction
Previous studies have mostly created training examples from stand-alone images which is essentially a reconstruction task, leading the resulting models to favor replication, often resulting in ’copy-and-paste’ artifacts [6]. To address this issue, our enhanced approach involves extracting subjects from a single video frame and using another frame from the same sequence as a reference for the truth. This technique effectively separates the personalized references from the target images. Due to possible variations in subjects between frames, we use a specialized subject-matching algorithm to ensure accurate matching. We provide a detailed description of this data processing pipeline in Section A.
@@ 3.3 Grounding Resampler
Different from text embedding, image embedding generally contains more information and is sparser, making projection into the condition space challenging. Leveraging embeddings from all image patches primarily control the condition, but the pooled output from the image encoder tends to omit many details. Different from [1] and [38], we propose the integration of a grounding resampler that functions as an alternative form of image projector. Utilizing a set of learnable tokens, the grounding resampler queries and distills pertinent information from the image features. Specifically, with an image embedding $f_{i}$ and a learnable query $f_{q}$, the resampler comprises several attention layers:
$\text{RSAttn}=\text{Softmax}\left(\frac{\mathbf{Q}\left(f_{q}\right)\mathbf{K}\left(\left[f_{i},f_{q}\right]\right)}{\sqrt{d}}\right)\mathbf{V}\left(\left[f_{i},f_{q}\right]\right)$ (3)
where $[f_{i},f_{q}]$ denotes the concatenation of the image embedding $f_{i}$ and the learnable query $f_{q}$. The architecture incorporates fully connected feedforward networks (FFNs), analogous to those utilized in standard vision transformers [8].
As detailed in 3.2, we can obtain entities of multiple referenced subjects and their target area boxes in the generated image. We present to initialize the query $f_{q}$ with grounding tokens $f_{g}$ derived from text embedding of entities and Fourier embedding of boxes. Entities are related to the semantic information of images and boxes indicate the areas where the subjects are supposed to be. It would be helpful for the resampler to extract the image features appropriately and the cross-attention layers to condition the generation finely. To prevent the model from becoming dependent on the grounding tokens during inference, we randomly replace them with the original learnable queries in the training. For the input of n subjects, the projection processes of different subject images do not affect each other. The resulting n queries will be concatenated and input into the subsequent model as $\mathbf{c_{i}}$ with $N=n*n_{t}$ tokens, where $n_{t}$ is the token quantity per subject.

### v2

# MS-Diffusion: Multi-subject Zero-shot Image Personalization with Layout Guidance

@@ Abstract
Recent advancements in text-to-image generation models have dramatically enhanced the generation of photorealistic images from textual prompts, leading to an increased interest in personalized text-to-image applications, particularly in multi-subject scenarios. However, these advances are hindered by two main challenges: firstly, the need to accurately maintain the details of each referenced subject in accordance with the textual descriptions; and secondly, the difficulty in achieving a cohesive representation of multiple subjects in a single image without introducing inconsistencies. To address these concerns, our research introduces the MS-Diffusion framework for layout-guided zero-shot image personalization with multi-subjects. This innovative approach integrates grounding tokens with the feature resampler to maintain detail fidelity among subjects. With the layout guidance, MS-Diffusion further improves the cross-attention to adapt to the multi-subject inputs, ensuring that each subject condition acts on specific areas. The proposed multi-subject cross-attention orchestrates harmonious inter-subject compositions while preserving the control of texts. Comprehensive quantitative and qualitative experiments affirm that this method surpasses existing models in both image and text fidelity, promoting the development of personalized text-to-image generation. The project page is https://MS-Diffusion.github.io.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion methodologies (Rombach et al., 2022; Saharia et al., 2022; Betker et al., 2023) have propelled the field to new heights, realizing unprecedented levels of photorealism while demonstrating a refined ability to conform to textual prompts. These achievements have spurred the development of a broad spectrum of applications, most prominently in the domain of personalized T2I (P-T2I) models, which are tasked with the complex undertaking of assimilating and regenerating novel visual concepts or subjects across diverse contexts with a heightened demand for conceptual and compositional fidelity. Despite fine-tuning-based techniques such as DreamBooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2023) yield results with considerable accuracy, they necessitate extensive resources for tuning individual instances and for the storage of multiple models, which renders them less feasible for widespread application. To circumvent these resource-intensive requirements, fine-tuning-free alternatives have come to the fore.
Single-subject driven personalization methods, IP-Adapter (Ye et al., 2023) and ELITE (Wei et al., 2023) for instance, introduce a specialized cross-attention mechanism that distinctly processes text and image features, thereby affording the possibility to employ reference images directly as visual prompts within the model. Furthermore, recent works have employed multi-subject driven customization methodologies to concatenate visual concepts with textual prompts, offering a glimpse of the potential in techniques like SSR-Encoder (Zhang et al., 2024), $\lambda$-ECLIPSE (Patel et al., 2024), Emu2 (Sun et al., 2023), and KOSMOS-G (Pan et al., 2023). These models harness identity data and amalgamate it with text via cross-attention, exhibiting proficiency in adjusting textures. Nevertheless, above zero-shot personalization methods encounter limitations, notably in adapting a pressing question pertains to the congruence of granular details between the subject depicted in the synthesized imagery and its corresponding subjects, along with the degree of semblance between the content of the generated image and associated textual descriptions. The challenge is further amplified in scenarios requiring the personalization of multiple subjects. Especially, the challenge of ensuring harmonious representation when multiple subjects are incorporated—specifically, the elucidation of whether the resultant image manifests any discordant elements or deleterious interactions in accordance with textual directives and multi-subject referential controls. As illustrated in Figure 2, multi-subject personalization methods frequently incur notable detail inaccuracies in a fine-tuning-free framework and often lead to subject neglect, subject overcontrol, and subject-subject conflict issue within the generated images.
@@
Textual Inversion (Gal et al., 2023)
DreamBooth (Ruiz et al., 2023)
@@
BLIP-Diffusion (Li et al., 2023a)
IP-Adapter (Ye et al., 2023)
@@
Kosmos-G (Pan et al., 2023)
$\lambda$-ECLIPSE (Patel et al., 2024)
SSR-Encoder (Zhang et al., 2024)
@@ 2.1 Text-to-image Generation
Text-to-image generative models (Saharia et al., 2022; Bao et al., 2023; Esser et al., 2024; Podell et al., 2023) are capable of producing high-quality images using user-provided text prompts. In recent times, diffusion-based models have shown strong performance in text-to-image tasks. Stable Diffusion (Rombach et al., 2022) proposes conducting the diffusion process in latent space rather than pixel space, which reduces the sampling steps without compromising image quality. Kandinsky (Razzhigaev et al., 2023) takes both text embedding and image embedding as conditions to generate images more controllablely. DALLE-3 (Betker et al., 2023) recaptions the training data pairs and utilizes T5-XXL (Chung et al., 2022) as the text encoder to strengthen the prompt-following ability. StableCascade (Pernias et al., 2023) presents a cascaded architecture to leverage outputs of front stages as priors, further reducing the latent space. PixArt-$\alpha$ (Chen et al., 2023a) also employs a large T5 text encoder and replaces the original U-Net backbone with a transformer (Peebles & Xie, 2023). These models focus on the basic text-to-image ability and cannot handle the situation when users provide specific subjects.
To finely control text-to-image generation, some diffusion models support users in providing layout guidance. Layout Diffusion (Zheng et al., 2023) and GLIGEN (Li et al., 2023c) input positions and labels of bounding boxes into the diffusion model and train it to learn the layout information. DenseDiffusion (Kim et al., 2023) develops a training-free method and modulates the attention maps in the inference phase. Instance Diffusion (Wang et al., 2024b) and MIGC (Zhou et al., 2024) extend the layout-conditioned diffusion to the instance level, enabling the model to generate multiple objects with precise quantities. While layout-guided diffusion models have robust controllability, they cannot reference specific concepts, which is emphasized in personalized text-to-image generation.
@@ 2.2 Text-to-image Personalization
Text-to-image personalization (Han et al., 2023; Qiu et al., 2023; Chen et al., 2024; Shi et al., 2024; Hu et al., 2024) has attracted much attention in the community for its powerful referencing capability of both text and image prompts. Textual inversion (Gal et al., 2023) and DreamBooth (Ruiz et al., 2023) utilize an identifier in the text to bind the visual concept through fine-tuning. IP-Adapter (Ye et al., 2023) proposes a zero-shot personalized model by projecting the image embedding to the cross-attention layers. InstanceID (Wang et al., 2024a) develops an approach for identity personalization, replacing the image encoder with a face encoder and employing ControlNet (Zhang et al., 2023) to integrate the face landmarks. To narrow the gap between image and text prompts, Kosmos-G (Pan et al., 2023) and $\lambda$-ECLIPSE (Patel et al., 2024) conduct multi-modal training to unite the inputs by text-image interleaving. SSR-Encoder (Zhang et al., 2024) design a query network to extract a single subject from images with multiple subjects for personalization.
Though past research in this field has significantly enhanced the ability to reference single subjects, few studies have explored zero-shot multi-subject personalized models. Moreover, existing related works struggle to address conflicts in personalized generation with multiple subjects and generate bad results, which is precisely the focus of our work, to discuss and resolve these issues.
Figure 3: The overall pipeline of MS-Diffusion. It introduces two pivotal enhancements to the model: the grounding resampler and multi-subject cross-attention mechanisms. Firstly, the grounding resampler adeptly assimilates visual information, correlating it with specific entities and spatial constraints. Subsequently, the cross-attention mechanism facilitates precise interactions between the image condition and the diffusion latent within the multi-subject attention layers. Throughout the training phase, all components of the pre-existing diffusion model remain frozen.
@@ 3.1 Preliminaries
As a widely used diffusion model, Stable Diffusion (SD) (Rombach et al., 2022) conducts the diffusion process in the latent space. Given an image and a text prompt, SD encodes them into latent code $\mathbf{z}$ and condition embedding $\mathbf{c_{t}}$ utilizing VAE (van den Oord et al., 2017) and CLIP (Radford et al., 2021) text encoder, respectively. In zero-shot image personalization architectures like IP-Adapter (Ye et al., 2023), images can also be considered a condition of the diffusion model. Specifically, a subject image is encoded to image embeddings by an image encoder and then projected into the original condition space of the diffusion model denoted as $\mathbf{c}_{i}$. For a timestep $t$ which is uniformly sampled from a fixed range, the model $\theta$ predicts the noise $\bm{\epsilon}_{\theta}$ and is optimized through the objective:
$\mathcal{L}_{IP}=\mathbb{E}_{\mathbf{z},\mathbf{c},\epsilon,t}\left[\left\|{\epsilon-\bm{\epsilon}}_{\theta}\left(\mathbf{z}_{t}\mid\mathbf{c_{t}},\mathbf{c_{i}},t\right)\right\|_{2}^{2}\right]$ (1)
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In this work, we employ $\mathcal{L}_{IP}$ with SDXL (Podell et al., 2023) as the pre-trained model, which contains two CLIP text encoders and additional condition inputs besides $\mathbf{c}$ and $t$, omitted for brevity.
@@
In IP-Adapter Ye et al. (2023), both $\mathbf{c}_{i}$ and $\mathbf{c}_{t}$ are integrated into the U-Net backbone through cross-attention layers:
$\text{Attn}\left(\mathbf{Q},\mathbf{K}_{i},\mathbf{K}_{t},\mathbf{V}_{i},\mathbf{V}_{t}\right)=\gamma\cdot\mathbf{z}_{img}+\mathbf{z}_{txt}=\gamma\cdot\text{Softmax}\left(\frac{\mathbf{Q}{\mathbf{K}_{i}^{\top}}}{\sqrt{d}}\right)\mathbf{V}_{i}+\text{Softmax}\left(\frac{\mathbf{Q}{\mathbf{K}_{t}^{\top}}}{\sqrt{d}}\right)\mathbf{V}_{t}$ (2)
where $\mathbf{Q}=\mathbf{z}W_{q}$, $\mathbf{K_{i}}=\mathbf{c}_{i}W_{k}^{i}$, $\mathbf{K_{t}}=\mathbf{c}_{t}W_{k}^{t}$, $\mathbf{V_{i}}=\mathbf{c}_{i}W_{v}^{i}$, $\mathbf{V_{t}}=\mathbf{c}_{t}W_{v}^{t}$, and $W_{q}$, $W_{k}$, $W_{v}$ are corresponding projection weight matrices, and $d$ represents the dimensionality of the key vectors. Note that the key and value matrix of $\mathbf{c}_{i}$ and $\mathbf{c}_{t}$ are independent of each other to decouple conditions of different modalities. Previous studies (Hertz et al., 2023; Tang et al., 2023) have found that attention maps $\mathbf{A}=\text{Softmax}\left(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d}\right)$ can reflect the attribution relation between generated images and conditions, which means that they determine the effect of condition controls.
3.2 Discussion on Multi-subject Image Personalization
A widely used method for achieving multi-subject image personalization involves training individual models for each subject, followed by their integration. Tuning-based methods (Gu et al., 2023; Avrahami et al., 2023; Liu et al., 2023b; Kumari et al., 2023) with improvements on visual conflicts can produce impressive multi-subject personalized images. However, zero-shot methods eliminate the need for individual subject tuning or the merging of different combinations, significantly reducing costs and enhancing the practicality of multi-subject personalization. This makes research into zero-shot multi-subject image personalization both necessary and promising.
Most of the relevant studies (Ma et al., 2024; Gu et al., 2023; Liu et al., 2023b; Xiao et al., 2024) focus on mitigating visual conflicts in text cross-attentions of the base model. While modifying text cross-attention can be effective, it presents certain limitations. First, adjustments to text cross-attention can directly impact the control over text conditions. Second, text cross-attention does not directly dictate the areas of influence for image conditions; rather, it exerts an indirect influence on image conditions by shaping the image layout generated by the diffusion model. This indirect control may result in low performance and increased uncertainty.
@@ 3.3 Data Construction
Previous studies have mostly created training examples from stand-alone images which is essentially a reconstruction task, leading the resulting models to favor replication, often resulting in ’copy-and-paste’ artifacts (Chen et al., 2023b). To address this issue, our enhanced approach involves extracting subjects from a single video frame and using another frame from the same sequence as a reference for the truth. This technique effectively separates the personalized references from the target images. Due to possible variations in subjects between frames, we use a specialized subject-matching algorithm to ensure accurate matching. We provide a detailed description of this data processing pipeline in Section A.
@@ 3.4 Grounding Resampler
Different from text embedding, image embedding generally contains more information and is sparser, making projection into the condition space challenging. Leveraging embeddings from all image patches primarily control the condition, but the pooled output from the image encoder tends to omit many details. Different from Flamingo (Alayrac et al., 2022) and IP-Adapter (Ye et al., 2023), we propose the integration of a grounding resampler that functions as an alternative form of image projector. Utilizing a set of learnable tokens, a resampler queries and distills pertinent information from the image features. Specifically, with an image embedding $f_{i}$ and a learnable query $f_{q}$, the resampler comprises several attention layers:
$\text{RSAttn}=\text{Softmax}\left(\frac{\mathbf{Q}\left(f_{q}\right){\mathbf{K}^{\top}}\left(\left[f_{i},f_{q}\right]\right)}{\sqrt{d}}\right)\mathbf{V}\left(\left[f_{i},f_{q}\right]\right)$ (3)
where $[f_{i},f_{q}]$ denotes the concatenation of the image embedding $f_{i}$ and the learnable query $f_{q}$. The architecture incorporates fully connected feedforward networks (FFNs), analogous to those utilized in standard vision transformers (Dosovitskiy et al., 2021).
As detailed in 3.3, we can obtain entities of multiple referenced subjects and their target area boxes in the generated image. We present to initialize the query $f_{q}$ with grounding tokens $f_{g}$ derived from text embedding of entities and Fourier embedding of boxes. Entities are related to the semantic information of images and boxes indicate the areas where the subjects are supposed to be. It would be helpful for the resampler to extract the image features appropriately and the cross-attention layers to condition the generation finely. To prevent the model from becoming dependent on the grounding tokens during inference, we randomly replace them with the original learnable queries in the training. For the input of n subjects, the projection processes of different subject images do not affect each other. The resulting n queries will be concatenated and input into the subsequent model as $\mathbf{c_{i}}$ with $N=n*n_{t}$ tokens, where $n_{t}$ is the token quantity per subject.